OpenStudy (anonymous):

PreCalculus help please! How do I transform this equation from polynomial form to vertex form by completing the square?! f(x)=x^2-8x=21

OpenStudy (*insert name here*):

You got the first step down, now do: \[\LARGE (\frac{b}{2})^2\]

OpenStudy (anonymous):

see your question please.

OpenStudy (anonymous):

Okay, so I get 16?

OpenStudy (anonymous):

if the eq is \[x ^{2}+ax=b,then add \left( \frac{ a }{ 2 } \right)^{2} \to both sides.\]

OpenStudy (*insert name here*):

I'm gonna wait and see what @surjithayer has to say

OpenStudy (*insert name here*):

So once you get 16, just like surj is showing, you add it to both sides so: \[\LARGE (x^2-8x+16)=21+16\] Now put it in: \[\LARGE (x \pm b)^2=c\]

OpenStudy (anonymous):

\[\left( x-4 \right)^{2}=37\]

OpenStudy (*insert name here*):

@surjithayer maybe we should let her do the work >.>

OpenStudy (anonymous):

Alright! Thank you guys, I think I get it now. I appreciate it! And I'm going over the work anyway @*Insert Name Here* , so don't worry. Thanks again! ♥

OpenStudy (*insert name here*):

Alright, good luck!